OpenStudy (anonymous):

convert to cylindrical coordinates. 3x+y-4z=12

OpenStudy (anonymous):

this is what i think should happen, but i'm not sure i'm allowed to do it... square both sides to get \[9x^2+y^2+16z^2 = 144\] then \[(9+16)r^2 = 144\] ??!!?

OpenStudy (amistre64):

thats not the square of both sides; thats just squaring each term that changes the value of the equation itself; ruins the rule

OpenStudy (anonymous):

oh..... :(

OpenStudy (anonymous):

k lemme try changing everything individually, one sec

OpenStudy (amistre64):

for instance: 5 + 3 = 8 ; square the terms like you did 25 + 9 = 64 is not right

OpenStudy (anonymous):

LOL, okay good point.. mashathomas.com/blush.jpg

OpenStudy (anonymous):

k it's actually easy lol.. \[3rcos \theta + rsin \theta -4z = 12\]

OpenStudy (anonymous):

^.^

OpenStudy (amistre64):

3x+y-4z=12 this forms a plane tilted in space; the z value is actually a slope defined in terms of the tilt

OpenStudy (anonymous):

then what if i have \[x^2-4z^2+y^2 = 0\]

OpenStudy (amistre64):